OpenStudy (anonymous):

the hypotenuse of a triangle is 13 less than twice one leg. If the other leg is 20cm, find the length of the other sides of the triangle.

OpenStudy (amistre64):



OpenStudy (amistre64):

now we pythag it to calibrate for the unknowns

OpenStudy (amistre64):

20^2 + L^2 = (2L-13)^2

OpenStudy (amistre64):

and by other leg im assuming its not the given leg

OpenStudy (amistre64):

400 + L^2 = 4L^2 -52L +196 0 = 3L^2 -52L -204 L = \(\cfrac{52\pm\sqrt{52^2-4(-204)(3)}}{2(3)}\) unless i seen it amis

OpenStudy (amistre64):

that makes it a wierd number; but about 20.63 for L

OpenStudy (amistre64):

which most likely means my brain lied to me along the way :)

OpenStudy (amistre64):

maybe i shoulda rounded up; 21^2 + 20^2 = (2(21)-13)^2 441+400=(42-13)^2 841=(29)^2 841=841 .... that did it